### v1

# Patch-Level Tokenization with CNN Encoders and Attention for Improved Transformer Time-Series Forecasting

Abstract
Transformer-based models have shown strong performance in time-series forecasting by leveraging self-attention to model long-range temporal dependencies. However, their effectiveness depends critically on the quality and structure of input representations derived from raw multivariate time-series data. This work proposes a two-stage forecasting framework that explicitly separates local temporal representation learning from global dependency modelling. In the first stage, a convolutional neural network (CNN) operates on fixed-length temporal patches to extract short-range temporal dynamics and non-linear feature interactions, producing compact patch-level token embeddings. Token-level self-attention is subsequently applied during representation learning to refine these embeddings by enabling interactions across temporal patches. In the second stage, a Transformer encoder processes the resulting token sequence to model inter-patch temporal dependencies and generate per-patch forecasts. Experiments conducted on synthetic multivariate time-series data with controlled static and dynamic factors demonstrate that the proposed patch-based tokenization strategy achieves competitive forecasting performance compared to convolutional and patch-based Transformer baselines. The results highlight the importance of structured temporal representations and show that decoupling local temporal encoding from global attention-based modelling yields more effective and stable time-series forecasting.
I Introduction
Accurately forecasting multivariate time-series data remains a challenging problem due to the complex, noisy, and often non-stationary nature of many real-world temporal processes. Such data are influenced by a combination of short-term fluctuations, longer-term trends, and interactions among multiple variables, resulting in non-linearity, heterogeneous feature dependencies, and evolving temporal dynamics. These characteristics complicate the design of effective forecasting models across a wide range of application domains.
Classical approaches to time-series forecasting, including autoregressive and moving-average models [12, 3], rely on assumptions of linearity and stationarity that are frequently violated in practice. As a result, deep learning methods have been increasingly adopted for modelling complex temporal behaviour. Recurrent neural networks (RNNs) [2], particularly Long Short-Term Memory (LSTM) [7] and Gated Recurrent Unit (GRU) [13] architectures, are among the most widely used models due to their ability to process sequential data and capture temporal dependencies. However, recurrent models summarize historical information through compressed hidden states, which can limit their ability to retain long-range dependencies and may lead to unstable training behaviour when long input sequences are considered.
Transformer-based architectures [14] have recently emerged as a powerful alternative for time-series forecasting. By employing self-attention mechanisms, Transformers enable direct interactions between all elements of a sequence, removing the reliance on recursive state propagation. This allows more effective modelling of long-range temporal dependencies and improved parallelization during training. Several studies have reported promising results using Transformer models for time-series forecasting [9, 15]. Nevertheless, Transformer performance is highly sensitive to the structure and quality of the input representations provided to the attention mechanism.
Unlike text or image data, multivariate time-series do not possess a natural tokenization scheme. When Transformers are applied directly to raw sequences, the model must simultaneously learn low-level feature extraction, noise suppression, and high-level temporal relationships. In many time-series settings, predictive information is concentrated in localized temporal patterns, such as short-term dynamics and volatility-like behaviour driven by recent observations [5, 4]. Attempting to capture both local temporal structure and long-range dependencies within a single attention-based model can therefore lead to inefficient use of model capacity and suboptimal representation learning [1, 11].
Motivated by recent advances in representation learning, which suggest that explicitly separating local pattern extraction from global dependency modelling can improve learning efficiency and performance [8], this work adopts a patch-based representation learning strategy for multivariate time-series forecasting. The proposed framework operates on fixed-length temporal patches, treating them as the fundamental units of representation. In the first stage, a convolutional neural network (CNN) encodes non-overlapping temporal patches to extract short-range temporal dynamics and non-linear feature interactions. Token-level self-attention is subsequently applied during representation learning to refine these patch embeddings by enabling interactions across temporal segments. In the second stage, a Transformer encoder processes the resulting sequence of patch-level tokens to model inter-patch temporal dependencies and generate forecasts.
The proposed method is evaluated on synthetic multivariate time-series datasets designed to reflect key structural properties commonly observed in real-world temporal data, including non-linear dynamics, interactions between static and dynamic factors, and stochastic noise. This controlled experimental setting enables systematic evaluation of the proposed representation learning strategy without introducing domain-specific confounding effects. Performance is compared against convolutional and patch-based Transformer baselines, including Temporal Convolutional Networks (TCN) and PatchTST, and results demonstrate that patch-level temporal tokenization consistently achieves competitive forecasting accuracy on unseen data.
II Methodology
II-A Proposed method
Let $\mathbf{X}=\{x_{t}\}_{t=1}^{T}$, where $x_{t}\in\mathbb{R}^{F}$, denote a multivariate time series. Instead of applying a forecasting model directly to $\mathbf{X}$, the proposed method first transforms the input into a sequence of learned temporal tokens. The time series is divided into non-overlapping temporal patches of length $P$, producing $K=\lfloor T/P\rfloor$ segments. Each segment is processed independently using a shared convolutional neural network, which extracts local temporal patterns and non-linear feature interactions. The convolutional outputs are aggregated using temporal pooling and projected into a $D$-dimensional embedding space, yielding a token sequence
$\mathbf{Z}\in\mathbb{R}^{K\times D}$. To incorporate contextual information across segments, multi-head self-attention is applied to the token sequence, enabling interactions across temporal patches during representation learning. The refined token sequence is subsequently used as input to a Transformer encoder, which models inter-patch temporal dependencies and produces patch-level forecasts. This design separates local representation learning from global temporal modelling, leading to more stable and effective forecasting.
Figure 1: Overview of the proposed CNN–Transformer architecture for multivariate time-series forecasting.
II-B CNN Patch Encoder + Attention
The convolutional neural network is employed as a patch encoder used to transform raw multivariate time-series input into optimized temporal token embeddings. A convolutional neural network is employed in the first stage due to its strong inductive bias toward extracting localized and stable patterns, making it well suited for modelling static or slowly varying feature interactions in multivariate time-series data. Subsequent token-level self-attention is applied to the patch embeddings produced by the convolutional encoder, enabling interactions across temporal patches during token refinement, extending beyond the local convolutional receptive field [6]. Given an input sequence $\mathbf{X}\in\mathbb{R}^{T\times F}$, where $T$ is the sequence length and $F$ is the number of input features, the sequence is first partitioned into non-overlapping fixed length temporal patches $P$. This results in a sequence of $K=\lfloor T/P\rfloor$ patches, each representing a localized temporal segment of the original series.
Each patch $\mathbf{X}_{k}\in\mathbb{R}^{P\times F}$ is processed independently using a shared one-dimensional convolutional encoder. Convolutions are applied along the temporal dimension, treating the feature dimension as channels, allowing the network to capture short-term temporal patterns and non-linear interactions among features within each patch. Dense connections are employed within the convolutional blocks to preserve intermediate representations and promote feature reuse, enabling both low- and higher-level temporal features to contribute to the final representation.
Following convolutional processing, learned attention-weighted pooling is applied across the temporal dimension of each patch to obtain a fixed-length vector representation. This pooled representation is projected into a $D$-dimensional embedding space, producing an initial token embedding for each patch. To incorporate contextual information across patches, multi-head self-attention is applied to the sequence of token embeddings. This allows each token to attend to others, refining the representations by integrating information from neighbouring and distant temporal segments. The output of this module is a sequence of optimized temporal token embeddings, which serve as structured inputs for subsequent Transformer-based modelling.
II-C Specialised Transformer
The Transformer stage receives as input a sequence of token embeddings $\mathbf{X}\in\mathbb{R}^{B\times K\times D_{\text{in}}}$, where $B$ is the batch size, $K$ is the number of temporal tokens, and $D_{in}$ is the embedding dimension produced by the CNN encoder. If the embedding dimension of the input differs from the model dimension $d_{\text{model}}$, a linear projection is first applied to map the input tokens to a common latent space.
$\mathbf{X}^{\prime}=\mathbf{X}\mathbf{W}_{p}+b_{p}$
$\mathbf{W}_{p}\in\mathbb{R}^{D_{in}\times d_{model}}$
To preserve temporal ordering, learned positional embeddings are added to the projected tokens. For each temporal patch $k\in\{1,\dots,K\}$, a positional vector $\mathbf{p}_{k}\in\mathbb{R}^{d_{\text{model}}}$ is retrieved from an embedding table and combined with the token representation,
$\mathbf{\tilde{x}}_{k}=\mathbf{x}^{\prime}_{k}+\mathbf{p}_{k}$
Dropout is applied to the position-enhanced sequence to regularize training.
The resulting sequence is processed by a stack of Transformer encoder layers operating in a pre-normalization configuration. Each layer applies multi-head self-attention followed by a position-wise feed-forward network, with residual connections and layer normalization to stabilize optimization. Self-attention enables each token to aggregate information from all other tokens in the sequence, allowing the model to capture long-range temporal dependencies.
Finally, a linear output head is applied to the Transformer outputs to produce forecasts at a fixed patch-level horizon. Given a forecast horizon $h>0$, the model predicts future patch-level targets according to
$\hat{y}_{k+h}=\mathbf{w}^{\top}\mathbf{h}_{k}+b,$
where $\mathbf{h}_{k}$ denotes the Transformer output corresponding to temporal patch $k$. The Transformer is trained by minimizing a regression loss between the predicted sequence and the corresponding ground-truth patch-level targets, while the patch encoder remains fixed.
Ground-truth targets are aligned to the patch structure by aggregating the original time-series values within each patch using a fixed reduction operation.
II-D Prediction Mechanism
After training, the model generates forecasts by mapping encoded temporal representations to future patch-level target values. Given an input sequence of length $T$, the Transformer encoder produces a sequence of hidden states $\{\mathbf{h}_{k}\}_{k=1}^{K}$, where each hidden state $\mathbf{h}_{k}\in\mathbb{R}^{D}$ corresponds to a specific temporal patch and encodes information aggregated from all patches in the input sequence through self-attention.
For a fixed forecast horizon $h$, predictions are generated using hidden states up to patch $k$ to forecast the target at patch $k+h$. A linear output layer is applied to each hidden state to obtain the predicted sequence,
$\hat{y}_{k+h}=\mathbf{w}^{\top}\mathbf{h}_{k}+b,\quad k=1,\dots,K-h,$
where $\mathbf{w}$ and $b$ are learned parameters. The model is trained to minimize a regression loss between the predicted patch-level forecasts $\{\hat{y}_{k+h}\}$ and the corresponding ground-truth patch-level targets. At inference time, the same horizon-aligned mapping is applied to previously unseen input sequences, producing a sequence of one-step-ahead patch-level forecasts aligned with the temporal patch structure of the input.
III Experimental setup
III-A Datasets
III-A1 Preparation
The experiments are conducted on a synthetic multivariate time-series dataset designed to provide controlled interactions between dynamic and static features [8]. Each sample consists of a time series of length $T$, where each observation $x_{t}\in\mathbb{R}^{6}$ contains two dynamic features and four static features.
The first dynamic feature $d_{t}^{(1)}$ is generated as a non-linear function of time,
$d_{t}^{(1)}=f(t)+\epsilon_{t}^{(1)},$
where $f(t)$ combines polynomial, exponential, and sinusoidal components, and $\epsilon_{t}^{(1)}\sim\mathcal{N}(0,\sigma_{1}^{2})$ denotes Gaussian noise.
The second dynamic feature $d_{t}^{(2)}$ is generated as a non-linear function of the first dynamic feature together with additional time-dependent components,
$d_{t}^{(2)}=\rho\,g(d_{t}^{(1)})+h(t)+p(t)+\epsilon_{t}^{(2)},$
where $g(\cdot)$ denotes a non-linear transformation involving linear and logarithmic terms, $h(t)$ denotes a trend component, $p(t)$ denotes a periodic component, $\rho$ controls the dependency strength, and $\epsilon_{t}^{(2)}\sim\mathcal{N}(0,\sigma_{2}^{2})$.
Four static features $\{s_{1},s_{2},s_{3},s_{4}\}$ are sampled once per sequence and remain constant across all time steps. These static features are repeated along the temporal dimension and concatenated with the dynamic features to form the input at each time step,
$x_{t}=[d_{t}^{(1)},d_{t}^{(2)},s_{1},s_{2},s_{3},s_{4}].$
The target sequence $\mathbf{Y}=\{y_{t}\}_{t=1}^{T}$ is generated as a non-linear combination of dynamic and static features,
$y_{t}=\alpha_{1}d_{t}^{(1)}s_{1}+\alpha_{2}\frac{35-s_{2}}{s_{2}-3}+\alpha_{3}d_{t}^{(2)}s_{4}+\epsilon_{t}^{(y)},$
where $\alpha_{1}=0.045$, $\alpha_{2}=0.38$, $\alpha_{3}=0.07$, and $\epsilon_{t}^{(y)}\sim\mathcal{N}(0,\sigma_{y}^{2})$.
The resulting dataset consists of input tensors $\mathbf{X}\in\mathbb{R}^{N\times T\times 6}$ and target tensors $\mathbf{Y}\in\mathbb{R}^{N\times T}$.
III-A2 Seeding
To ensure a fair and reproducible evaluation, two independent datasets were generated using random seeds 42 and 101 for training and testing, respectively. Both datasets follow identical data generation and preprocessing pipelines. All proposed and baseline models are trained exclusively on the training dataset and evaluated on the held-out test dataset.
III-A3 Illustrative Sample
Table I presents an illustrative example from the synthetic dataset, showing the first few time steps of a single multivariate time-series sample. Each observation consists of two dynamic features and four static features, where the static features remain constant across time steps.
TABLE I: Illustrative example of a single synthetic multivariate time-series sample (first 6 time steps). Static features remain constant across time.
$t$ $d_{t}^{(1)}$ $d_{t}^{(2)}$ $s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$
1 -5.36 0.36 2 16.43 3 3
2 -4.52 6.41 2 16.43 3 3
3 -2.87 3.73 2 16.43 3 3
4 -1.38 2.48 2 16.43 3 3
5 -2.02 6.28 2 16.43 3 3
6 -1.76 2.45 2 16.43 3 3
The dynamic features exhibit non-linear and noisy temporal behaviour, including trend and oscillatory components, while the static features act as sequence-level conditioning variables that modulate the target generation process. As shown in the example, static features remain fixed across time steps within a sequence, whereas the dynamic features vary over time. This structure reflects scenarios commonly encountered in financial and economic time-series data, where asset-specific or regime-specific factors remain constant over short horizons while market signals evolve temporally.
III-B Compared Models
The proposed method is evaluated against representative baseline models that capture different design paradigms commonly used in time-series forecasting. In particular, we compare against a convolutional sequence model and a patch-based Transformer model, which serve as strong and widely adopted baselines for modelling local and global temporal dependencies.
Temporal Convolutional Network (TCN)
Temporal Convolutional Networks [1] employ stacks of one-dimensional convolutional layers with increasing receptive fields to model temporal dependencies in sequential data. By using dilated convolutions, TCNs are able to capture long-range temporal patterns while maintaining a fully convolutional structure and avoiding recurrent connections. This design enables stable training and parallel computation, making TCNs a strong baseline for time-series forecasting tasks. In our experiments, the TCN operates directly on the raw multivariate input sequence and produces patch-aligned forecasts using a comparable training protocol.
PatchTST
PatchTST [11] is a recent Transformer-based time-series forecasting model that introduces a patch-level tokenization strategy prior to attention modelling. Instead of applying self-attention directly to individual time steps, the input sequence is partitioned into fixed-length temporal patches, which are linearly embedded and treated as tokens for a Transformer encoder. This approach reduces sequence length, improves computational efficiency, and encourages the model to focus on patch-level temporal structure. PatchTST serves as a strong baseline for evaluating the effectiveness of patch-based representations in Transformer architectures.
All baseline models are trained and evaluated under the same experimental conditions as the proposed method, using identical data splits, input features, and evaluation metrics, to ensure a fair and consistent comparison.
III-C Implementation details
The proposed model and other compared models have been implemented in python with the PyTorch library; architecture and experimentation is publicly available [10]. All models are trained using the AdamW optimizer.
III-D Hyperparameter Settings
TABLE II: Key hyperparameter settings used in all experiments.
Parameter Value
Number of input features 6
Sequence length ($T$) 30
Patch size ($P$) 8
Number of patches ($\lfloor T/P\rfloor$) 3
Forecasting granularity Patch-level
TABLE III: Hyperparameter settings for the proposed model and baseline methods.
Model Architecture Layers Attention Heads Learning Rate Epochs Batch Size
Proposed CNN Patch Encoder $1\times 10^{-3}$ 2000 32
Transformer Encoder 4 16 $1\times 10^{-3}$ 300 32
TCN Temporal Conv. Network 4 $1\times 10^{-3}$ 300 32
PatchTST Patch-based Transformer 4 16 $1\times 10^{-3}$ 300 32
Tables II and III summarize the global and model-specific hyperparameter settings used in the experiments
IV Results
The forecasting performance of the evaluated models is assessed using Mean Squared Error (MSE) and Mean Absolute Error (MAE), defined as
$MSE=\frac{1}{NT}\sum^{N}_{i=1}\sum^{T}_{t=1}(y_{i,t}-\hat{y}_{i,t})^{2},$
$MAE=\frac{1}{NT}\sum^{N}_{i=1}\sum^{T}_{t=1}\lvert y_{i,t}-\hat{y}_{i,t}\rvert$,
where $N$ denotes the number of samples, $T$ the sequence length, $y_{i,t}$ the ground-truth target value at time step $t$ for sample $i$, and $\hat{y}_{i,t}$ the corresponding predicted value. Both metrics are computed at the patch level and averaged over all predicted patches.
Figure 2: Comparison of forecasting performance across models using MSE and MAE.
Figure 2 compares the forecasting performance of the proposed model with the TCN and PatchTST baselines. All models achieve low error values, indicating that the synthetic dataset is well suited for evaluating patch-based forecasting approaches. Among the evaluated methods, PatchTST attains the lowest error values, achieving an MSE of 0.028 and an MAE of 0.131. The proposed model achieves an MSE of 0.093 and an MAE of 0.232, while the TCN baseline records an MSE of 0.038 and an MAE of 0.153.
The observed differences in performance can be interpreted in light of the architectural characteristics of the models. PatchTST applies self-attention directly to patch-level representations that are learned end-to-end for the forecasting objective, which may be advantageous in the considered synthetic setting where target values depend primarily on contemporaneous dynamic features and static covariates. This direct optimization of patch embeddings enables efficient modelling of non-linear feature interactions.
The TCN baseline achieves competitive performance, highlighting the effectiveness of convolutional architectures with large receptive fields for capturing localized temporal structure and reducing noise. This suggests that a substantial portion of the predictive signal in the dataset is contained within short-range temporal patterns that can be effectively modelled without explicit attention mechanisms.
The proposed two-stage framework demonstrates stable and consistent performance across both error metrics. By decoupling local temporal representation learning from global dependency modelling, the approach produces structured patch-level tokens that remain informative for downstream forecasting, even though the representation learning and forecasting stages are trained separately. While the proposed method does not outperform PatchTST on this benchmark, it achieves comparable performance to strong convolutional and Transformer-based baselines, supporting the effectiveness of patch-level temporal tokenization for multivariate time-series forecasting.
V Conclusion
This study investigates a multi-stage forecasting architecture that explicitly separates patch-level temporal tokenization from global sequence modelling in multivariate time-series forecasting. By first learning compact patch-level token embeddings through a convolutional encoder and subsequently refining these representations using token-level self-attention prior to Transformer-based modelling, the proposed framework captures localized temporal structure while enabling global dependency modelling at a coarser temporal resolution.
Empirical results on a controlled synthetic dataset designed to reflect key structural properties of financial time series demonstrate that the proposed approach achieves competitive forecasting performance relative to strong convolutional and patch-based Transformer baselines. While a direct patch-based Transformer model achieves the lowest prediction errors on this benchmark, the proposed method performs comparably across both MSE and MAE, indicating that decoupling local temporal representation learning from global attention-based modelling does not degrade predictive accuracy.
These findings suggest that offloading local temporal representation learning from the Transformer may provide a structured and stable alternative to end-to-end attention-based modelling, particularly in settings where localized temporal patterns and feature interactions play an important role. The modular nature of the proposed framework also facilitates representation reuse and architectural flexibility, which may be advantageous in more complex or heterogeneous forecasting scenarios.
While the current implementation employs a convolutional patch encoder for token generation, future work may explore alternative patch-level encoders, including recurrent or hybrid architectures, as well as jointly trained or task-aware representation learning strategies. Although the evaluation is conducted on synthetic data, the results highlight the potential of structured patch-level tokenization for multivariate time-series forecasting, with the understanding that performance on real-world financial data may depend on domain-specific characteristics and market conditions.
Acknowledgment
The authors acknowledge the support provided by VIT-AP University during the course of this research.