### v3

# Patch-Level Tokenization with CNN Encoders and Attention for Improved Transformer Time-Series Forecasting

Abstract
Transformer-based models have shown strong performance in time-series forecasting by leveraging self-attention to model long-range temporal dependencies. However, their effectiveness depends critically on the quality and structure of input representations derived from raw multivariate time-series data, particularly as sequence length and data scale increase. This paper proposes a two-stage forecasting framework that explicitly separates local temporal representation learning from global dependency modelling. In the proposed approach, a convolutional neural network operates on fixed-length temporal patches to extract short-range temporal dynamics and non-linear feature interactions, producing compact patch-level token embeddings. Token-level self-attention is applied during representation learning to refine these embeddings, after which a Transformer encoder models inter-patch temporal dependencies to generate forecasts. The method is evaluated on a synthetic multivariate time-series dataset with controlled static and dynamic factors, using an extended sequence length and a larger number of samples. Experimental results demonstrate that the proposed framework consistently outperforms a convolutional baseline under increased temporal context and remains competitive with a strong patch-based Transformer model. These findings indicate that structured patch-level tokenization provides a scalable and effective representation for multivariate time-series forecasting, particularly when longer input sequences are considered.
I Introduction
Accurately forecasting multivariate time-series data remains a challenging problem due to the complex, noisy, and often non-stationary nature of many real-world temporal processes. Such time-series data are shaped by a mix of short-lived fluctuations, longer-term trends, and interactions between multiple variables. Together, these factors introduce non-linear behaviour, uneven feature dependencies, and temporal patterns that change over time. As a result, designing forecasting models that perform reliably across different application domains remains a challenging task. Traditional time-series forecasting methods, such as autoregressive and moving-average models [12, 3], are built on assumptions of linearity and stationarity. In many real-world settings, these assumptions do not hold up, which limits the effectiveness of such an approach. This has led to growing interest in deep learning techniques for modelling complex temporal behaviour. Recurrent neural networks (RNNs) [2] have been among the most widely adopted of these methods, as they are specifically designed to process sequential data, particularly Long Short-Term Memory (LSTM) [7] and Gated Recurrent Unit (GRU) [13] architectures, are among the most widely adopted models due to their ability to process sequential data and capture local temporal dependencies. However, recurrent models summarize historical information through compressed hidden states, which can limit their ability to retain long-range dependencies and may lead to unstable training behaviour when long input sequences are considered.
Transformer-based architectures [16] have recently emerged as a powerful alternative for time-series forecasting. By employing self-attention mechanisms [15], Transformers enable direct interactions between all elements of a sequence, removing the reliance on recursive state propagation. This allows more effective modelling of long-range temporal dependencies and improved parallelization during training. Several studies have reported promising results using Transformer models for time-series forecasting [9, 17]. Nevertheless, Transformer performance is highly sensitive to the structure and quality of the input representations provided to the attention mechanism.
Unlike text or image data, multivariate time-series do not possess a natural tokenization scheme. When Transformers are applied directly to raw sequences, the model must simultaneously learn low-level feature extraction, noise suppression, and high-level temporal relationships. In many time-series settings, predictive information is concentrated in localized temporal patterns, such as short-term dynamics and volatility-like behaviour driven by recent observations [5, 4]. Attempting to capture both local temporal structure and long-range dependencies within a single attention-based model can therefore lead to inefficient use of model capacity and suboptimal representation learning [1, 11].
Motivated by recent advances in representation learning, which suggest that explicitly separating local pattern extraction from global dependency modelling can improve learning efficiency and performance [8], this work adopts a patch-based representation learning strategy for multivariate time-series forecasting. The proposed framework operates on fixed-length temporal patches, treating them as the fundamental units of representation. In the first stage, a convolutional neural network (CNN) encodes non-overlapping temporal patches to extract short-range temporal dynamics and non-linear feature interactions. Token-level self-attention is subsequently applied during representation learning to refine these patch embeddings by enabling interactions across temporal segments. In the second stage, a Transformer encoder processes the resulting sequence of patch-level tokens to model inter-patch temporal dependencies and generate forecasts.
The proposed architecture is initially evaluated on a constructed synthetic dataset designed to reflect key properties and structural dependencies as is observed in real world data, including static and dynamic features as well as stochastic noise. Performance is compared against convolutional and patch-based Transformer baselines: Temporal Convolutional Networks (TCN) [1] and PatchTST [11], and results demonstrate that patch-level temporal tokenization consistently achieves competitive forecasting accuracy on unseen data. Model is then evaluated on a real dataset to show architecture capability in real scenarios and not just tailored situations.
The remainder of this paper is organized as follows. Section II describes the proposed patch-based forecasting framework, detailing the CNN-based patch encoder, token-level self-attention mechanism, and the specialized Transformer architecture used for global temporal modelling. Section III presents the experimental setup, including datasets used and how synthetic dataset was generated, baseline models, implementation details, and hyperparameter settings. Section IV reports the empirical results and provides a comparative analysis of the proposed method against the baseline approaches. Section V concludes the paper and draws subjective conclusions for the proposed model. Section VI discusses the limitations of current architecture, how it can be improved and scope for future works.
II Methodology
II-A Proposed method
Let $\mathbf{X}=\{x_{t}\}_{t=1}^{T}$, where $x_{t}\in\mathbb{R}^{F}$, denote a multivariate time series of length $T$. Rather than applying a forecasting model directly to the raw input sequence, the proposed approach first transforms $\mathbf{X}$ into a sequence of learned temporal tokens. The time series is partitioned into non-overlapping temporal patches of fixed length $P$, resulting in $K=\lfloor T/P\rfloor$ segments. Each patch is processed independently by a shared convolutional neural network, which captures short-range temporal patterns and non-linear interactions among features within the patch. The convolutional representations are subsequently aggregated along the temporal dimension and projected into a $D$-dimensional embedding space, producing a sequence of patch-level tokens $\mathbf{Z}\in\mathbb{R}^{K\times D}$. To enable information exchange across temporal segments, multi-head self-attention is applied to the token sequence during representation learning, allowing dependencies between patches to be explicitly modelled. The refined token sequence is subsequently used as input to a Transformer encoder, which models inter-patch temporal dependencies and produces patch-level forecasts. This design separates local representation learning from global temporal modelling, leading to more stable and effective forecasting.
Figure 1: Overview of the proposed CNN–Transformer architecture for multivariate time-series forecasting.
II-B CNN Patch Encoder and Attention
The convolutional neural network is employed as a patch encoder used to transform raw multivariate time-series input into optimized temporal token embeddings. A convolutional neural network is employed in the first stage due to its strong inductive bias toward extracting localized and stable patterns, making it well suited for modelling static or slowly varying feature interactions in multivariate time-series data. Subsequent token-level self-attention is applied to the patch embeddings produced by the convolutional encoder, enabling interactions across temporal patches during token refinement, extending beyond the local convolutional receptive field [6]. Given an input sequence $\mathbf{X}\in\mathbb{R}^{T\times F}$, where $T$ is the sequence length and $F$ is the number of input features, the sequence is first partitioned into non-overlapping fixed length temporal patches $P$. This results in a sequence of $K=\lfloor T/P\rfloor$ patches, each representing a localized temporal segment of the original series.
Each patch $\mathbf{X}_{k}\in\mathbb{R}^{P\times F}$ is processed independently using a shared one-dimensional convolutional encoder. Convolutions are applied along the temporal dimension, treating the feature dimension as channels, allowing the network to capture short-term temporal patterns and non-linear interactions among features within each patch. Dense connections are applied within the convolutional blocks of the encoder to preserve intermediate representations and promote future use, enabling the contribution of both low and higher-level temporal features in final representation. Following the convolutional processing, a learned attention weighted pooling is applied across the temporal dimension of each patch to obtain a fixed length vector representation. This pooled representation is then projected into a $D$-dimensional embedding space, thus producing an initial token embedding for each patch. To incorporate contextual information across patches, multi-head self-attention is applied to the sequence of token embeddings. This allows each token to attend to others, refining the representations by integrating information from neighbouring and distant temporal segments. The output of this module is a sequence of optimized temporal token embeddings, which serve as structured inputs for subsequent Transformer-based modelling.
II-C Specialised Transformer
The Transformer stage receives as input a sequence of token embeddings $\mathbf{X}\in\mathbb{R}^{B\times K\times D_{\text{in}}}$, where $B$ is the batch size, $K$ is the number of temporal tokens, and $D_{in}$ is the embedding dimension produced by the CNN encoder. If the embedding dimension of the input differs from the model dimension $d_{\text{model}}$, a linear projection is first applied to map the input tokens to a common latent space.
$\mathbf{X}^{\prime}=\mathbf{X}\mathbf{W}_{p}+b_{p}$
$\mathbf{W}_{p}\in\mathbb{R}^{D_{in}\times d_{model}}$
To preserve temporal ordering, learned positional embeddings are added to the projected tokens. For each temporal patch $k\in\{1,\dots,K\}$, a positional vector $\mathbf{p}_{k}\in\mathbb{R}^{d_{\text{model}}}$ is retrieved from an embedding table and combined with the token representation,
$\mathbf{\tilde{x}}_{k}=\mathbf{x}^{\prime}_{k}+\mathbf{p}_{k}$
Dropout is applied to the position-enhanced sequence to regularize training.
The resulting sequence is processed by a stack of Transformer encoder layers operating in a pre-normalization configuration. Each layer applies multi-head self-attention followed by a position-wise feed-forward network, with residual connections and layer normalization to stabilize optimization. Self-attention enables each token to aggregate information from all other tokens in the sequence, allowing the model to capture long-range temporal dependencies.
Finally, a linear output head is applied to the Transformer outputs to produce forecasts at a fixed patch-level horizon. Given a forecast horizon $h>0$, the model predicts future patch-level targets according to
$\hat{y}_{k+h}=\mathbf{w}^{\top}\mathbf{h}_{k}+b,$
where $\mathbf{h}_{k}$ denotes the Transformer output corresponding to temporal patch $k$. The Transformer is trained by minimizing a regression loss between the predicted sequence and the corresponding ground-truth patch-level targets, while the patch encoder remains fixed.
Ground-truth targets are aligned to the patch structure by aggregating the original time-series values within each patch using a fixed reduction operation.
II-D Prediction Mechanism
After training, the model generates forecasts by mapping encoded temporal representations to future patch-level target values. Given an input sequence of length $T$, the Transformer encoder produces a sequence of hidden states $\{\mathbf{h}_{k}\}_{k=1}^{K}$, where each hidden state $\mathbf{h}_{k}\in\mathbb{R}^{D}$ corresponds to a specific temporal patch and encodes information aggregated from all patches in the input sequence through self-attention.
For a fixed forecast horizon $h$, predictions are generated using hidden states up to patch $k$ to forecast the target at patch $k+h$. A linear output layer is applied to each hidden state to obtain the predicted sequence,
$\hat{y}_{k+h}=\mathbf{w}^{\top}\mathbf{h}_{k}+b,\quad k=1,\dots,K-h,$
where $\mathbf{w}$ and $b$ are learned parameters. The model is trained to minimize a regression loss between the predicted patch-level forecasts $\{\hat{y}_{k+h}\}$ and the corresponding ground-truth patch-level targets. At inference time, the same horizon-aligned mapping is applied to previously unseen input sequences, producing a sequence of one-step-ahead patch-level forecasts aligned with the temporal patch structure of the input.
III Experimental setup
III-A Datasets
III-A1 Preparation
The experiments are conducted on a synthetic multivariate time-series dataset designed to provide controlled interactions between dynamic and static features [8]. Each sample consists of a time series of length $T$, where each observation $x_{t}\in\mathbb{R}^{6}$ contains two dynamic features and four static features.
The first dynamic feature $d_{t}^{(1)}$ is generated as a non-linear function of time,
$d_{t}^{(1)}=f(t)+\epsilon_{t}^{(1)},$
where $f(t)$ combines polynomial, exponential, and sinusoidal components, and $\epsilon_{t}^{(1)}\sim\mathcal{N}(0,\sigma_{1}^{2})$ denotes Gaussian noise.
The second dynamic feature $d_{t}^{(2)}$ is generated as a non-linear function of the first dynamic feature together with additional time-dependent components,
$d_{t}^{(2)}=\rho\,g(d_{t}^{(1)})+h(t)+p(t)+\epsilon_{t}^{(2)},$
where $g(\cdot)$ denotes a fixed non-linear transformation composed of linear and logarithmic terms, $h(t)$ denotes a trend component, $p(t)$ denotes a periodic component, $\rho$ controls the dependency strength, and $\epsilon_{t}^{(2)}\sim\mathcal{N}(0,\sigma_{2}^{2})$.
Four static features $\{s_{1},s_{2},s_{3},s_{4}\}$ are sampled once per sequence and remain constant across all time steps. These static features are repeated along the temporal dimension and concatenated with the dynamic features to form the input at each time step,
$x_{t}=[d_{t}^{(1)},d_{t}^{(2)},s_{1},s_{2},s_{3},s_{4}].$
The target sequence $\mathbf{Y}=\{y_{t}\}_{t=1}^{T}$ is generated as a non-linear combination of dynamic and static features,
$y_{t}=\alpha_{1}d_{t}^{(1)}s_{1}+\alpha_{2}\frac{35-s_{2}}{s_{2}-3}+\alpha_{3}d_{t}^{(2)}s_{4}+\epsilon_{t}^{(y)},$
where $\alpha_{1}=0.045$, $\alpha_{2}=0.38$, $\alpha_{3}=0.07$, and $\epsilon_{t}^{(y)}\sim\mathcal{N}(0,\sigma_{y}^{2})$. The noise variances $\sigma_{1}^{2}$, $\sigma_{2}^{2}$, and $\sigma_{y}^{2}$ are fixed constants controlling the noise level in the synthetic data generation process.
The resulting dataset consists of input tensors $\mathbf{X}\in\mathbb{R}^{N\times T\times 6}$ and target tensors $\mathbf{Y}\in\mathbb{R}^{N\times T}$.
III-A2 Seeding
To ensure a fair and reproducible evaluation, two independent datasets were generated using random seeds 42 and 101 for training and testing, respectively. Both datasets follow identical data generation and preprocessing pipelines. All proposed and baseline models are trained exclusively on the training dataset and evaluated on the held-out test dataset.
III-A3 Illustrative Sample
Table I presents an illustrative example from the synthetic dataset, showing the first few time steps of a single multivariate time-series sample. Each observation consists of two dynamic features and four static features, where the static features remain constant across time steps.
TABLE I: Illustrative example of a single synthetic multivariate time-series sample (first 6 time steps). Static features remain constant across time.
$t$ $d_{t}^{(1)}$ $d_{t}^{(2)}$ $s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$
1 -5.36 0.36 2 16.43 3 3
2 -4.52 6.41 2 16.43 3 3
3 -2.87 3.73 2 16.43 3 3
4 -1.38 2.48 2 16.43 3 3
5 -2.02 6.28 2 16.43 3 3
6 -1.76 2.45 2 16.43 3 3
The dynamic features exhibit non-linear and noisy temporal behaviour, including trend and oscillatory components, while the static features act as sequence-level conditioning variables that modulate the target generation process. As shown in the dataset snippet, static features are fixed across time-steps within a sequence, the dynamic features, however, vary over time sequence. This structure reflects scenarios commonly encountered in financial and economic time-series data, where asset-specific or regime-specific factors remain constant over short horizons while market signals evolve temporally.
III-B Compared Models
The proposed method is evaluated against representative baseline models that capture different design paradigms commonly used in time-series forecasting. In particular, we compare against a convolutional sequence model and a patch-based Transformer model, which serve as strong and widely adopted baselines for modelling local and global temporal dependencies.
Temporal Convolutional Network (TCN)
Temporal Convolutional Networks (TCNs) [1] model temporal dependencies using stacks of one-dimensional dilated convolutions, allowing large receptive fields without recurrent connections. Their fully convolutional structure enables stable training and parallel computation, making them a widely used baseline for time-series forecasting. In our experiments, the TCN operates directly on the raw multivariate input sequence and produces patch-aligned forecasts under a comparable training protocol.
PatchTST
PatchTST [11] is a Transformer-based forecasting model that applies self-attention to fixed-length temporal patches rather than individual time steps. By treating patches as tokens, PatchTST reduces sequence length and focuses attention on patch-level temporal structure. It serves as a strong baseline for evaluating patch-based Transformer approaches in time-series forecasting.
All baseline models are trained and evaluated under the same experimental conditions as the proposed method, using identical data splits, input features, and evaluation metrics, to ensure a fair and consistent comparison.
III-C Implementation details
The proposed model and other compared models have been implemented in python with the library; architecture and experimentation is publicly available [10]. All models are trained using the AdamW optimizer.
III-D Hyperparameter Settings
Tables II and III summarize the global and model-specific hyperparameter settings used in the experiments.
TABLE II: Key hyperparameter settings used in all experiments.
Parameter Value
Number of input features 6
Sequence length ($T$) 160
Patch size ($P$) 8
Number of patches ($\lfloor T/P\rfloor$) 20
Forecasting granularity Patch-level
TABLE III: Hyperparameter settings for the proposed model and baseline methods.
Model Architecture Layers Attention Heads Learning Rate Epochs Batch Size
Proposed CNN Patch Encoder $1\times 10^{-3}$ 2000 32
Transformer Encoder 4 16 $1\times 10^{-3}$ 300 32
TCN Temporal Conv. Network 4 $1\times 10^{-3}$ 300 32
PatchTST Patch-based Transformer 4 16 $1\times 10^{-3}$ 300 32
IV Results
The forecasting performance of the evaluated models is assessed using Mean Squared Error (MSE) and Mean Absolute Error (MAE), defined as
$MSE=\frac{1}{NT}\sum^{N}_{i=1}\sum^{T}_{t=1}(y_{i,t}-\hat{y}_{i,t})^{2},$
$MAE=\frac{1}{NT}\sum^{N}_{i=1}\sum^{T}_{t=1}\lvert y_{i,t}-\hat{y}_{i,t}\rvert$,
where $N$ denotes the number of samples, $T$ the sequence length, $y_{i,t}$ the ground-truth target value at time step $t$ for sample $i$, and $\hat{y}_{i,t}$ the corresponding predicted value. Both metrics are computed at the patch level and averaged over all predicted patches.
IV-A Synthetic Dataset
All results reported in this section are obtained using a dataset of N=10, [^0] N=10{,}000 samples with sequence length $T=160$.
Figure 2 compares the forecasting performance of the proposed model with the TCN and PatchTST baselines. All models achieve stable performance under the larger dataset size and extended sequence length, indicating that the synthetic dataset remains suitable for evaluating patch-based forecasting approaches at scale. Among the evaluated methods, PatchTST attains the lowest error values, achieving an MSE of 0.055 and an MAE of 0.185. The proposed model achieves an MSE of 0.074 and an MAE of 0.194, while the TCN baseline records an MSE of 0.199 and an MAE of 0.306.
Figure 2: Comparison of forecasting performance across models using MSE and MAE.
The differences in forecasting performance can be understood by considering the architectural design of each model. PatchTST applies self-attention directly to patch-level representations that are learned end-to-end for the forecasting task, which appears beneficial in the synthetic setting considered here, especially when longer input sequences are used. By jointly learning patch embeddings and attention weights, the model is able to capture non-linear feature interactions in an efficient and task-aligned manner.
In contrast, the TCN baseline exhibits higher error under extended sequence lengths, indicating that relying solely on convolutional receptive fields may limit the ability to model longer-range temporal dependencies. While TCNs effectively model localized temporal structure, their performance degrades relative to attention-based approaches as the temporal context increases.
The proposed two-stage framework demonstrates consistently improved performance over the convolutional baseline and remains competitive with the patch-based Transformer model. By decoupling local temporal representation learning from global dependency modelling, the approach produces structured patch-level tokens that remain informative for downstream forecasting, even though representation learning and forecasting are performed in separate stages. These results indicate that structured patch-level tokenization continues to be effective as sequence length and dataset scale increase.
IV-B Real-World Evaluation: Electricity Load Forecasting
To assess the generalization capability of the proposed framework beyond controlled synthetic settings, we further evaluate the model on the Electricity Load Diagrams 2011–2014 dataset from the UCI Machine Learning Repository [14]. This dataset consists of electricity consumption measurements recorded at 15-minute intervals for a large number of individual customers over multiple years and is widely used as a benchmark for long-horizon time-series forecasting.
Following standard practice, the data are normalized on a per-meter basis and segmented into sliding windows of fixed length. Each input sample consists of a multivariate sequence of length $T=160$, with the forecasting target defined as the patch-level average load of a selected meter. The same patch size and forecasting horizon used in the synthetic experiments are retained to ensure a consistent evaluation protocol across datasets. Notably, the convolutional patch encoder trained on synthetic data is reused without architectural modification, reflecting the length-agnostic nature of the learned local temporal representations.
On the Electricity Load Diagrams dataset, the proposed model achieves a Mean Squared Error (MSE) of 0.0136 and a Mean Absolute Error (MAE) of 0.0624 on the test split. These results indicate accurate short-horizon forecasting in a real-world setting characterized by strong periodic structure, noise, and long-range temporal dependencies. The relatively low MAE value, computed on normalized data, suggests that the model is able to capture both fine-grained load variations and broader consumption trends.
The observed performance can be linked to the architectural separation between local and global modelling stages. Electricity consumption data exhibit smooth intra-day dynamics alongside recurring daily and weekly patterns. The convolutional patch encoder processes short temporal segments, reducing local noise and extracting stable short-range features, while the Transformer focuses on modelling dependencies across patches at a coarser temporal scale. This division of roles enables the attention mechanism to operate on more structured and compact inputs, which appears particularly beneficial when longer temporal contexts are considered.
Overall, these results demonstrate that the proposed patch-based tokenization strategy generalizes beyond synthetic data and remains effective on a large-scale real-world benchmark. Although a detailed comparison with specialized electricity forecasting models is outside the scope of this study, the findings provide empirical support for the benefits of decoupling local temporal representation learning from global attention-based modelling in practical multivariate time-series forecasting tasks.
V Conclusion
This study has aimed to present a multi-stage forecasting model that explicitly separates patch-level tokenization from global sequence modelling in MTSF. By compacting patches through a convolutional encoder and then learning them, refining them through self-attention prior to Transformer modelling, the proposed framework captures localised temporal structure while enabling effective modelling of longer range dependencies.
Empirical results on a large synthetic dataset with extended sequence length show that the proposed approach consistently outperforms a purely convolutional baseline and achieves competitive performance relative to a strong patch-based Transformer model. Additional evaluation on a real-world electricity load forecasting benchmark demonstrates that the learned patch-level representations transfer effectively beyond controlled synthetic data. While the proposed method does not outperform PatchTST in all settings, it exhibits improved robustness over convolutional baselines and stable performance under increased temporal context.
VI Limitations and Future Work
A limitation of this study is that a substantial portion of the analysis relies on synthetic data, which, while enabling controlled investigation of model behaviour, may not fully capture the complexity and regime variability present in real-world time-series. Although additional evaluation on electricity load data demonstrates promising generalization, further validation on diverse real-world benchmarks is required.
In addition, the convolutional patch encoder and Transformer components are trained in separate stages, which may restrict the extent to which task-specific information is incorporated into the learned representations. Future work will explore jointly trained architectures, alternative patch-level encoders, and adaptive tokenization strategies to further improve representation quality and forecasting performance.
Acknowledgment
The authors acknowledge the support provided by VIT-AP University during the course of this research.